OpenStudy (anonymous):

0.3x-0.2y=4 0.5x+0.5y=-15/19 solve by elimination method

OpenStudy (anonymous):

Lcm for -0.2 and 0.5 is 0.10

OpenStudy (anonymous):

someone correct me if i'm wrong. but i got x=820/19 (43.15789474)

OpenStudy (anonymous):

panther do you get at all what you are supposed to do here?

OpenStudy (anonymous):

im not sure i understand this one

OpenStudy (anonymous):

Solve by Elimination. In Both of (0.3x and 0.5x) look for a Least Common factor. and that is (0.15x) For (-0.2 and 0.5) the Least common factor is (0.10)

OpenStudy (anonymous):

ok then what would i do next

OpenStudy (anonymous):

pick one that satisfied you better. for instance .15x is Least common factor for 0.3 and 0.5. so we multiply (0.3x-0.2y=4)*0.5 so we end up with 0.15x (0.5x+0.5y=-15/19)*-0.15 so we end up with x variable Negative -x0.15

OpenStudy (anonymous):

ok so we would then eliminate x right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

make sure you make your calculations correctly

OpenStudy (anonymous):

this is when we would divide or am i getting ahead of myself